please help time is running out for me and I simply just dont know the anwser!
The sum of two numbers is 90. The second is 10 more than 4 times the first. What are the two numbers?
Is there anyone out there that can solve this problem?

Answer by ankor@dixie-net.com(12689)

Write a simple equation for each statement:
:
The sum of two numbers is 90.
x + y = 90
:
The second is 10 more than 4 times the first,
y = x + 10
What are the two numbers?
Since y = x+10 we can substitute it for y in the 1st equation:
x + (x+10) = 90
2x + 10 = 90
2x = 90 - 10
x = 80/2
x = 40
:
If x = 40 then obviously, y = 50, which is 10 more that x